# Permutation twisted cohomology, remixed

### v2

@@ 3. Chain complexes over the orbit category
Since effective endotrivials identify with representation spheres, they have additional structural properties, as representation spheres (and more generally, any $G$-CW-complex) produce free chain complexes over the orbit category $\Gamma_{G}$. We first review this notion, following [Yal17], then prove some technical results which will be critical for the sequel.
@@
Let $\Gamma_{G}$ denote the orbit category of $G$. The objects of $\Gamma_{G}$ are transitive $G$-sets $G/H$ for subgroups $H\leq G$, and the morphisms from $G/H$ to $G/K$ are $G$-set homomorphisms $G/H\to G/K$. In particular, $\operatorname{Hom}_{\Gamma_{G}}(H,K)$ is empty unless a conjugate of $H$ is a subgroup of $K$.
A $k\Gamma_{G}$-module $M$ is a contravariant functor from the category $\Gamma_{G}$ to the category of $k$-modules. By identifying $\operatorname{Aut}_{\Gamma_{G}}(G/H)$ with G//H{G}/\!\!/{H}, $M(H)$ has k[G//H]k[{G}/\!\!/{H}]-module structure. The category of finitely generated $k\Gamma_{G}$-modules, denoted $\operatorname{mod}(\Gamma_{G})$ is abelian, so the usual categorical concepts apply.
Given a $G$-set $X$, we define the $k\Gamma_{G}$-module $k[X^{?}]$ as the module with value at $G/H$ given by $k[X^{H}]$, with the obvious induced maps. A module over $\Gamma_{G}$ is free if it is isomorphic to a direct sum of modules of the form $k[(G/K)^{?}]$. Let $\operatorname{proj}(\Gamma_{G})$ denote the full subcategory of $\operatorname{mod}(\Gamma_{G})$ consisting of free $\Gamma_{G}$-modules (by the Yoneda lemma, every projective $\Gamma_{G}$-module is free, see [Yal17, Definition 3.1]).
@@
Let $X$ and $Y$ be two $G$-sets. Any $k\Gamma_{G}$-homomorphism $f\colon k[X^{?}]\to k[Y^{?}]$ satisfies the following property: given any $x\in X$ and subgroup $H\leq G$, if $x\in X^{H}$, then $f(x)\in k[Y^{H}]$. In particular, $\operatorname{Hom}_{\operatorname{mod}(\Gamma_{G})}(k[X^{?}],k[Y^{?}])$ has $k$-basis induced from the set $\operatorname{Hom}_{G-\operatorname{set}}(X,Y)$.
@@
As a result, each free $k\Gamma_{G}$-module has a stratification by fixed points. That is, given a free $k\Gamma_{G}$-module $kX$, we may write $kX=kG^{\oplus a_{1}}\oplus k[G/H_{2}]^{\oplus a_{2}}\cdots\oplus k[G/G]^{a_{n}}$ for some $a_{1},\dots,a_{n}\in\mathbb{N}$, where each direct summand $k[G/H_{i}]^{\oplus a_{i}}$ is a uniquely determined submodule of $kX$. Therefore, we may fix a “canonical permutation basis” in accordance with this stratification. We will make considerable use of this fact in the sequel.
Given a $G$-set $X$ and two subgroups $K\leq H$ of $G$, we have a canonical inclusion homomorphism (of $k[N_{G}(H)\cap N_{G}(K)]$-modules) $i\colon k[X^{H}]\to k[X^{K}]$ associated to the $G$-set homomorphism $G/K\to G/H,1K\mapsto 1H$. This comes associated with a (not necessarily unique) projection map $p\colon k[X^{K}]\to k[X^{H}]$ satisfying $p\circ i=\operatorname{id}$ and $p(m)=0$ if $m\in\Psi^{H}(k[X^{K}])=0$. After choosing a canonical permutation basis, we obtain a corresponding projection map associated to the basis.
For any subgroup $H\leq G$, we have an obvious functor $(H)\colon\operatorname{proj}(\Gamma_{G})\to\operatorname{perm}(k[{G}/\!\!/{H}])$ from the assignment $k[X^{?}]\to k[X^{H}]$. Modular fixed points behave as one would hope.
@@
$\operatorname{proj}(\Gamma_{G})$$\operatorname{perm}(kG)$$\operatorname{perm}(k[{G}/\!\!/{P}])$$\scriptstyle{-(1)}$$\scriptstyle{-(P)}$$\scriptstyle{\Psi^{P}}$
@@
Let $X$ be a $G$-CW-complex. The reduced chain complex $X$ over the orbit category is the functor $\tilde{C}_{\ast}(X^{?};k)$ from the orbit category $\Gamma_{G}$ to the category of chain complexes of $k$-modules. This gives rise to a chain complex of free (by [Yal17, Lemma 3.2]) $k\Gamma_{G}$-modules
$\tilde{C}_{\ast}(X^{?};k):=\cdots\to C_{i}(X^{?};k)\xrightarrow{d_{i}}C_{i-1}(X^{?};k)\to\cdots\to C_{0}(X^{?};k)\xrightarrow{\epsilon}\underline{k}\to 0,$
where $\underline{k}$ denotes the constant functor with values $k(H)=k$ for all subgroups $H\leq G$, and $\epsilon$ denotes the augmentation homomorphism (this is essentially the construction of reduced Bredon homology, see [Ill73]). We denote chain complexes of free $k\Gamma_{G}$-modules by $C^{?}$, and their evaluations at the subgroup $H\leq G$ by $C^{H}$ to inspire an aura of fixed points.
If $G$ is a finite $p$-group and $S$ is a representation sphere of $G$, it may be realized as a $G$-CW-complex, hence producing a chain complex of free $k\Gamma_{G}$-modules. Therefore for any subgroup $H\leq G$, $\tilde{C}_{\ast}(S^{H};k)$ is an effective endotrivial complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules. Conversely, after shifting, every effective endotrivial complex $C$ of $kG$-modules is homotopy equivalent to $\tilde{C}_{\ast}(S^{1};k)$ for some representation sphere $S$ of $G$ with dimension homomorphism $\dim(S)=h_{C}$ (see [tD87, Page 217, Theorem 5.16]), and hence up to homotopy arises from a chain complex of free $k\Gamma_{G}$-modules.
This proposition is critical for many of the constructions to follow: it allows us to perform lifts of $p$-local homomorphisms $k\to\Psi^{H}(C)$.
@@
Let $G$ be a finite $p$-group, let $k[X^{?}]$ and $k[Y^{?}]$ be two free $k\Gamma_{G}$-modules with $X$ transitive, let $f\colon k[X^{?}]\to k[Y^{?}]$ be a $k\Gamma_{G}$-module homomorphism, and let $H\leq G$ be a subgroup. Suppose $k[X^{H}]\neq 0$, and let $M^{\prime}$ denote the unique minimal nonzero submodule of $k[X^{H}]$. If $M^{\prime}\subseteq\ker(f^{H})$, then there exists a submodule $M^{?}\subseteq k[X^{?}]$, minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$, such that $M^{H}=M^{\prime}$ and for all subgroups $K\leq_{G}H$, $M^{K}\subseteq\ker(f^{K})$.
@@
We define $M^{?}\subseteq k[X^{?}]$ as follows: for $L\leq G$, $M^{L}\subset k[X^{L}]$ is the unique minimal submodule if $X$ has any $K$-fixed points, and the zero module if not. It is straightforward that, by construction, $M^{?}$ is minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$. Now by minimality of $M^{?}$, it suffices to show $f^{K}$ is not an injective k[G//K]k[{G}/\!\!/{K}]-module homomorphism for all $K\leq_{G}H$, but this follows directly from Proposition 3.3. ∎
@@
Proposition 3.7 asserts that we can construct “local” inclusions $k\to C[s]$ for chain complexes arising from chain complexes over the orbit category. This does not hold for arbitrary objects of $\mathscr{K}(G)$. For instance, the two-term complex $C=k\xrightarrow{\operatorname{coaug}}kG$ (with $k$ in homological degree 0) has no nonzero global homomorphism $k\to C$, but there exist nonzero local homomorphisms $k\to\Psi^{H}(C)$ (in fact isomorphisms) for every nontrivial subgroup $H\leq G$, since $\Psi^{H}(kG)=0$.
@@ 3.1. Removing contractible summands
We prove a couple technical results for finite $p$-groups that show we can remove contractible summands from chain complexes of free $k\Gamma_{G}$-modules if and only if we can for the corresponding chain complexes of permutation $kG$-modules (in particular, for effective endotrivials). Though most of these results are not strictly necessary for future sections, they provide elucidation and may be of independent interest.
@@
A chain complex of free $k\Gamma_{G}$ modules has in each homological degree a canonical permutation basis, and each differential respects the stabilizers grow condition. Moreover, it is easy to see that given a chain complex $C$ of permutation $kG$-modules, if one can choose a canonical permutation basis of $G$-sets in each degree such that the stabilizers grow condition holds with respect to this basis, then $C$ may be realized as the image of a chain complex of free $k\Gamma_{G}$-modules evaluated at the subgroup $1\leq G$. In this sense, the chain complexes of permutation $kG$-modules satisfying the stabilizers grow condition on differentials are exactly those which arise from chain complexes of free $k\Gamma_{G}$-modules.
If $C$ is a chain complex of permutation modules, we say $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules if there exists a chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C^{1}\cong C$ as chain complexes.
@@
It follows from Corollary 3.3 that there exists an index $a\in\{1,\dots,m\}$ such that the composition $p_{b}\circ d_{i}\circ i_{a}$ is an isomorphism, where $p_{b}$ denotes projection onto $kY_{b}$ and $i_{a}$ denotes inclusion into $kX_{a}$. In particular, $X_{a}$ and $Y_{b}$ are isomorphic $G$-sets. Note the choice of $a$ is not necessarily unique (for example, the two-term complex $k\oplus k\to k$ with differential $(\operatorname{id},\operatorname{id})$).
Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{a}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}^{\prime}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{b}\oplus\cdots\oplus kY_{n}$. We will now modify the canonical bases of $C_{i}$ and $C_{i-1}$ to construct the contractible chain complex $K$ that splits off of $C$. Choose any generator $t\in T$ of $T$, then we may write $t=p_{a}(t)+t^{\prime}$, with $t^{\prime}\in T_{1}\oplus\cdots\oplus\widehat{T}_{a}\oplus\cdots\oplus T_{m}$. The $kG$-modules $\langle p_{a}(t)\rangle$ and $\langle t^{\prime}\rangle$ are both isomorphic to $k$, i.e. are $G$-stable. We replace the $C_{i}$ basis elements $\{x_{1},\dots,x_{e}\}=X_{a}\subset kX_{a}\subseteq C_{i}$ with $\{x_{1}+t^{\prime},\dots,x_{e}+t^{\prime}\}=X_{a}^{\prime}\subset C_{i}$. $X_{a}^{\prime}$ is again a $G$-set, and we have $C_{i}=C_{i}^{\prime}\oplus kX^{\prime}_{a}$. Moreover, $T\subseteq kX_{a}^{\prime}$ by construction.
We replace the $C_{i-1}$ basis elements $\{y_{1},\dots,y_{f}\}=Y_{b}\subset kY_{b}\subseteq C_{i-1}$ with $\{d_{i}(x_{1}+t^{\prime}),\dots,d_{i}(x_{e}+t^{\prime})\}=X_{b}^{\prime}\subset C_{i-1}$. Again, $Y_{b}^{\prime}$ is a $G$-set. We claim that $p_{b}(X_{a}^{\prime})$ is a $k$-basis of $Y_{b}$. This follows because $\langle p_{b}\circ d_{i}(t^{\prime})\rangle$ is either 0 or the unique $k$-dimension one submodule of $kY_{b}$, and we chose $b$ such that both $p_{b}\circ d_{i}\circ i_{a}$ was an isomorphism and $p_{b}\circ d_{i}(T)$ was nonzero. Therefore $p_{b}\circ d_{i}(y_{j}+t^{\prime})$ has the same $G$-stabilizer as both $y_{j}$ and $y_{j}+t^{\prime}$. Hence $Y_{b}^{\prime}$ is isomorphic as a $G$-set to $Y_{b}$ and $X_{a}$, and it follows that $C_{i-1}=C^{\prime}_{i-1}\oplus kY^{\prime}_{b}$. Moreover, $d_{i}$ restricts to an isomorphism $kX_{a}^{\prime}\xrightarrow{\cong}kY_{b}^{\prime}$, and the result follows. ∎
The next proposition demonstrates that if a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules satisfies that $C^{1}$ has a contractible summand (as a chain complex of $kG$-modules), then $C^{?}$ itself has a contractible summand (as a chain complex of $k\Gamma_{G}$-modules).
@@
Let $G$ be a finite group and suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. If $C$ contains a contractible direct summand $K$, then there exists a direct sum decomposition $C=K\oplus D$ such that $D$ also arises from a chain complex $D^{?}$ of free $k\Gamma_{G}$-modules.
@@
To prove this, it suffices to find a decomposition of $C$ into direct summands $K,D$ as stated and show that each homological degree of $D$ has a canonical permutation basis satisfying the stabilizers grow condition. First, since $C$ has a contractible direct summand, there exists an integer $i\in\mathbb{Z}$ and $k$-dimension one submodule $T\subseteq C_{i}$ such that $T\not\in\ker(d_{i})$. Let $C_{i}=kX_{1}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}=kY_{1}\oplus\cdots\oplus kY_{n}$ be the direct sum decompositions into canonical bases, with each $X_{j},Y_{j}$ a transitive permutation module. After projecting $T$ onto each $kX_{j}$, it follows that there exists an index $l$ for which $d_{i}|_{X_{l}}$ is injective. Choose the index $l$ such that the stabilizer of $X_{l}$ is minimal with respect to subgroup inclusion.
@@
We modify the canonical bases of $C_{i}$ and $C_{i-1}$ as follows. First, it follows from Corollary 3.3 that there exists a (non-unique) index $h\in\{1,\dots,n\}$ such that $p_{h}\circ d_{i}\circ i_{l}$ is an isomorphism, where $i_{l}$ denotes inclusion into $kX_{l}$ and $p_{h}$ denotes projection onto $kY_{h}$. We replace the basis elements of $Y_{h}$ with the basis elements $Y_{h}^{\prime}=\{d_{i}(x)\mid x\in X_{l}\}$. Then $Y_{h}^{\prime}$ is a $G$-set isomorphic to $X_{l}$, and since the projection of $kY_{h}^{\prime}$ onto $kY_{h}$ is an isomorphism, this forms a new basis of $C_{i-1}$. Finally, we have $d_{i-1}(kY_{h}^{\prime})=0$ since $d_{i-1}\circ d_{i}=0$, so the resulting basis still respects the stabilizers grow property. Set $D_{i-1}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{h}\oplus\cdots\oplus kY_{n}$ and $K_{i-1}:=kY_{h}^{\prime}$.

### v3

@@ 3. Chain complexes over the orbit category
The connection between representation spheres, endotrivial complexes and Borel-Smith functions lets us discern that effective endotrivials have additional structure: they are induced from representation spheres via Bredon (alternatively called $G$-equivariant) homology. $G$-CW-complexes produce chain complexes of free modules over the orbit category $\Gamma_{G}$, which themselves induce particularly nice objects of $\mathscr{K}(G)$, well-behaved with respect to local-global constructions. We first review the orbit category, following [YAL17].
@@
Let $\Gamma_{G}$ denote the orbit category of $G$. The objects of $\Gamma_{G}$ are transitive $G$-sets $G/H$ for subgroups $H\leq G$, and the morphisms from $G/H$ to $G/K$ are $G$-set homomorphisms $G/H\to G/K$. In particular, $\operatorname{Hom}_{\Gamma_{G}}(H,K)$ is empty unless a conjugate of $H$ is a subgroup of $K$. We may write a subgroup $H\leq G$ to denote the object $G/H\in\Gamma_{G}$ for shorthand.
A $k\Gamma_{G}$-module $M$ is a contravariant functor from the category $\Gamma_{G}$ to the category of $k$-modules. By identifying $\operatorname{Aut}_{\Gamma_{G}}(G/H)$ with G//H{G}/\!\!/{H}, $M(H)$ has k[G//H]k[{G}/\!\!/{H}]-module structure. The category of finitely generated $k\Gamma_{G}$-modules, denoted $\operatorname{mod}(k\Gamma_{G})$ is abelian.
Given a $G$-set $X$, we define the $k\Gamma_{G}$-module $k[X^{?}]$ as the module with value at $G/H$ given by the subspace $k[X^{H}]$, with the obvious induced maps. A module over $\Gamma_{G}$ is free if it is isomorphic to a direct sum of modules of the form $k[(G/K)^{?}]$. Let $\operatorname{proj}(k\Gamma_{G})$ denote the full subcategory of $\operatorname{mod}(k\Gamma_{G})$ consisting of free $\Gamma_{G}$-modules (by the Yoneda lemma, every projective $\Gamma_{G}$-module is free, see [YAL17, Definition 3.1]).
@@
Let $X$ and $Y$ be two $G$-sets. Any $k\Gamma_{G}$-homomorphism $f\colon k[X^{?}]\to k[Y^{?}]$ satisfies the following property: given any $x\in X$ and subgroup $H\leq G$, if $x\in X^{H}$, then $f(x)\in k[Y^{H}]$. In particular, $\operatorname{Hom}_{\operatorname{mod}(k\Gamma_{G})}(k[X^{?}],k[Y^{?}])$ has $k$-basis induced from $\operatorname{Hom}_{G-\operatorname{set}}(X,Y)$.
@@
As a result, each free $k\Gamma_{G}$-module has a filtration by fixed points. Therefore, we may fix a “canonical permutation basis” in accordance with this stratification. We will make considerable use of this fact in the sequel.
Given a $G$-set $X$ and two subgroups $K\leq H$ of $G$, we have a canonical inclusion homomorphism of $k[N_{G}(H)\cap N_{G}(K)]$-modules $i\colon k[X^{H}]\to k[X^{K}]$ associated to the $G$-set homomorphism $G/K\to G/H,1K\mapsto 1H$. This comes associated with a (not necessarily unique) projection map $p\colon k[X^{K}]\to k[X^{H}]$ satisfying $p\circ i=\operatorname{id}$ and $p(m)=0$ if $m\in\Psi^{H}(k[X^{K}])=0$. After choosing a canonical permutation basis, we obtain an corresponding projection map associated to the basis.
For any subgroup $H\leq G$, we have an obvious functor $(H)\colon\operatorname{proj}(k\Gamma_{G})\to\operatorname{perm}(k[{G}/\!\!/{H}])$ from the assignment $k[X^{?}]\to k[X^{H}]$. Modular fixed points behave as one would hope.
@@
$\operatorname{proj}(k\Gamma_{G})$$\operatorname{perm}(kG)$$\operatorname{perm}(k[{G}/\!\!/{P}])$$\scriptstyle{-(1)}$$\scriptstyle{-(P)}$$\scriptstyle{\Psi^{P}}$
@@
We now explain how to obtain an endotrivial complex from a representation sphere, via a construction topologists will be familiar with.
@@
Let $X$ be a $G$-CW-complex. The reduced cellular chain complex of $X$ over the orbit category is the functor $\tilde{C}(X^{?}):=\tilde{C}(X^{?};k)$ from the orbit category $\Gamma_{G}$ to the category of chain complexes of $k$-modules. This gives rise to a chain complex of free $k\Gamma_{G}$-modules
$\tilde{C}(X^{?}):=\cdots\to C_{i}(X^{?})\xrightarrow{d_{i}}C_{i-1}(X^{?})\to\cdots\to C_{0}(X^{?})\xrightarrow{\epsilon}\underline{k}\to 0,$
where $\underline{k}$ denotes the constant functor with values $k(H)=k$ for all subgroups $H\leq G$, and $\epsilon$ denotes the augmentation homomorphism. By convention $\underline{k}$ is in homological degree -1. This is essentially the construction of reduced Bredon homology with coefficients in $k$, see [ILL73]. We denote chain complexes of free $k\Gamma_{G}$-modules by $C^{?}$, and their evaluations at the subgroup $H\leq G$ by $C^{H}$ to inspire an aura of fixed points.
Given any chain complex of free $k\Gamma_{G}$-modules, evaluation at $1$ produces a chain complex of $p$-permutation kGkG-modules. This assignment induces a tt-functor
$(-)^{1}\colon\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\mathscr{K}(G)$, and the analogous commutative diagram to 3.5 commutes.
Let $G$ be a $p$-group. If $S^{V}$ is a representation sphere of $G$, it may be realized as a $G$-CW-complex, hence producing a chain complex of free $k\Gamma_{G}$-modules. Therefore for any subgroup $H\leq G$, $\tilde{C}(S^{H})$ is the corresponding effective endotrivial complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules. It is clear there is an equality
$h_{C}(H)=\operatorname{H}_{i}(\tilde{C}(S^{H}))=\dim_{\mathbb{R}}V^{H}$
for all $H\leq G$.
Conversely, by identifying an endotrivial $C$ with its h-mark homomorphism $h_{C}$, Theorem 2.10 asserts $C$ is up to shift homotopy equivalent to $\tilde{C}(S^{1})$ for some representation sphere $S$ of $G$ with dimension homomorphism $\dim(S)=h_{C}$ by 3.5. Therefore, up to homotopy, every effective endotrivial arises from a chain complex of free $k\Gamma_{G}$-modules. We will show that one can remove “up to homotopy” in the sequel.
@@
Let $G$ be a finite $p$-group, let $k[X^{?}]$ and $k[Y^{?}]$ be two free $k\Gamma_{G}$-modules with $X$ transitive, let $f\colon k[X^{?}]\to k[Y^{?}]$ be a $k\Gamma_{G}$-module homomorphism, and let $H\leq G$ be a subgroup. Suppose $k[X^{H}]\neq 0$, and let $M^{\prime}$ denote the unique minimal nonzero submodule of $k[X^{H}]$. If $M^{\prime}\subseteq\ker(f^{H})$, then there exists a unique submodule $M^{?}\subseteq k[X^{?}]$, minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$, such that $M^{H}=M^{\prime}$ and for all subgroups $K\leq_{G}H$, $M^{K}\subseteq\ker(f^{K})$.
@@
We define $M^{?}\subseteq k[X^{?}]$ as follows: for $L\leq G$, $M^{L}\subset k[X^{L}]$ is the unique minimal submodule of $k$-dimension 1 if $X$ has any $L$-fixed points, and 0 if not. By construction, $M^{?}$ is minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$. Now by minimality of $M^{?}$, it suffices to show $f^{K}$ is not an injective k[G//K]k[{G}/\!\!/{K}]-module homomorphism for all $K\leq_{G}H$, and this follows from 3.3. ∎
@@
3.7 asserts that we can lift “$p$-local” morphisms $\hat{f}\colon k\to\Psi^{H}(C[s])$ to “global” morphisms $f\colon k\to C[s]$ satisfying $\Psi^{H}(f)=\hat{f}$ for chain complexes arising from chain complexes over the orbit category. This does not hold for arbitrary objects of $\mathscr{K}(G)$. For instance, the two-term complex $C=k\xrightarrow{\operatorname{coaug}}kG$ (with $k$ in homological degree 0) has no nonzero global homomorphism $k\to C$, but there exist nonzero local homomorphisms $k\to\Psi^{H}(C)$ (in fact isomorphisms) for every nontrivial subgroup $H\leq G$, since $\Psi^{H}(kG)=0$. These constructions make sense in $\mathscr{K}(G)$ rather than $\operatorname{K}_{b}(\operatorname{mod}(k\Gamma_{G}))$, as $\operatorname{Hom}_{\operatorname{mod}(k\Gamma_{G})}(\underline{k},k[X^{?}])=0$ for any nontrivial transitive $G$-set $X$.
@@ 3.1. Removing contractible summands
We prove a couple technical results for finite $p$-groups that show we can remove contractible summands from chain complexes of free $k\Gamma_{G}$-modules if and only if we can for the corresponding chain complexes of permutation $kG$-modules (in particular, for effective endotrivials).
@@
A chain complex of free $k\Gamma_{G}$ modules has in each homological degree a canonical permutation basis induced from the natural inclusions $X^{H}\hookrightarrow X^{K}$ for $K\leq_{G}H\leq G$, and each differential respects the stabilizers grow condition. Moreover, it is easy to see that given a chain complex $C$ of permutation $kG$-modules, if one can choose a canonical permutation basis of $G$-sets in each degree such that the stabilizers grow condition holds with respect to this basis, then $C$ may be realized as the image of a chain complex of free $k\Gamma_{G}$-modules evaluated at the subgroup $1\leq G$. In this sense, the chain complexes of permutation $kG$-modules satisfying the stabilizers grow condition on differentials with respect to some chosen permutation basis are exactly those which arise from chain complexes of free $k\Gamma_{G}$-modules.
If $C$ is a chain complex of permutation modules, we say $C$ arises from a chain complex of free $k\Gamma_{G}$-modules if $C$ is in the essential image of the “evaluation at G/1G/1” $t$-functor $(-)^{1}\colon\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\mathscr{K}(G)$.
@@
It follows from 3.3 that there exists an index $a\in\{1,\dots,m\}$ such that the composition $p_{b}\circ d_{i}\circ i_{a}$ is an isomorphism, where $p_{b}$ denotes projection onto $kY_{b}$ and $i_{a}$ denotes inclusion into $kX_{a}$. In particular, $X_{a}$ and $Y_{b}$ are isomorphic $G$-sets. Note the choice of $a$ is not necessarily unique (for example, the two-term complex $k\oplus k\to k$ with differential $(\operatorname{id},\operatorname{id})$).
Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{a}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}^{\prime}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{b}\oplus\cdots\oplus kY_{n}$. We will now modify the canonical bases of $C_{i}$ and $C_{i-1}$ to construct the contractible chain complex $K$ that splits off of $C$. Choose any generator $t\in T$ of $T$, then we may write $t=p_{a}(t)+t^{\prime}$, with $t^{\prime}\in T_{1}\oplus\cdots\oplus\widehat{T}_{a}\oplus\cdots\oplus T_{m}$. The $kG$-modules $\langle p_{a}(t)\rangle$ and $\langle t^{\prime}\rangle$ are both isomorphic to $k$, i.e., are $G$-stable. We replace the $C_{i}$ basis elements $\{x_{1},\dots,x_{e}\}=X_{a}\subset kX_{a}\subseteq C_{i}$ with $\{x_{1}+t^{\prime},\dots,x_{e}+t^{\prime}\}=X_{a}^{\prime}\subset C_{i}$. $X_{a}^{\prime}$ is again a $G$-set, and we have $C_{i}=C_{i}^{\prime}\oplus kX^{\prime}_{a}$. Moreover, $T\subseteq kX_{a}^{\prime}$ by construction.
We replace the $C_{i-1}$ basis elements $\{y_{1},\dots,y_{f}\}=Y_{b}\subset kY_{b}\subseteq C_{i-1}$ with $\{d_{i}(x_{1}+t^{\prime}),\dots,d_{i}(x_{e}+t^{\prime})\}=X_{b}^{\prime}\subset C_{i-1}$. Again, $Y_{b}^{\prime}$ is a $G$-set. We claim that $p_{b}(X_{a}^{\prime})$ is a $k$-basis of $Y_{b}$. This follows because $\langle p_{b}\circ d_{i}(t^{\prime})\rangle$ is either 0 or the unique $k$-dimension one submodule of $kY_{b}$, and we chose $b$ such that both $p_{b}\circ d_{i}\circ i_{a}$ was an isomorphism and $p_{b}\circ d_{i}(T)$ was nonzero. Therefore $p_{b}\circ d_{i}(y_{j}+t^{\prime})$ has the same $G$-stabilizer as both $y_{j}$ and $y_{j}+t^{\prime}$. Hence $Y_{b}^{\prime}$ is isomorphic as a $G$-set to $Y_{b}$ and $X_{a}$, and $C_{i-1}=C^{\prime}_{i-1}\oplus kY^{\prime}_{b}$. Moreover, $d_{i}$ restricts to an isomorphism $kX_{a}^{\prime}\xrightarrow{\cong}kY_{b}^{\prime}$, and the result follows. ∎
The next proposition demonstrates that if a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules satisfies that $C^{1}$ has a contractible summand (as a chain complex of $kG$-modules), then $C^{?}$ itself has a contractible summand (as a chain complex of $k\Gamma_{G}$-modules). This fact will be rather useful for studying effective endotrivials. It is unclear if a non-technical proof of this fact exists, but we suspect that for chain complexes arising from $G$-CW-complexes, it may be so.
@@
Let $G$ be a finite group and suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. If $C$ contains a contractible direct summand $K$ of permutation modules, then there exists a direct sum decomposition $C=K\oplus D$ such that $D$ also arises from a chain complex $D^{?}$ of free $k\Gamma_{G}$-modules.
@@
To prove this, it suffices to find a decomposition of $C$ into direct summands $K,D$ as stated and show that each homological degree of $D$ has a canonical permutation basis satisfying the stabilizers grow condition. If $C$ is indecomposable, there is nothing to show, so assume $C$ decomposes. First, since $C$ has a contractible direct summand, there exists an integer $i\in\mathbb{Z}$ and $k$-dimension one submodule $T\subseteq C_{i}$ such that $T\not\in\ker(d_{i})$. Let $C_{i}=kX_{1}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}=kY_{1}\oplus\cdots\oplus kY_{n}$ be the direct sum decompositions into canonical bases, with each $X_{j},Y_{j}$ a transitive permutation module. After projecting $T$ onto each $kX_{j}$, it follows that there exists an index $l$ for which $d_{i}|_{X_{l}}$ is injective. Choose an index $l$ such that the stabilizer of $X_{l}$ is minimal with respect to group order.
@@
We modify the canonical bases of $C_{i}$ and $C_{i-1}$ as follows. First, 3.3 implies there exists a (non-unique) index $h\in\{1,\dots,n\}$ such that $p_{h}\circ d_{i}\circ i_{l}$ is an isomorphism, where $i_{l}$ denotes inclusion into $kX_{l}$ and $p_{h}$ denotes projection onto $kY_{h}$. We replace the basis elements of $Y_{h}$ with the basis elements $Y_{h}^{\prime}=\{d_{i}(x)\mid x\in X_{l}\}$. Then $Y_{h}^{\prime}$ is a $G$-set isomorphic to $X_{l}$, and since the projection of $kY_{h}^{\prime}$ onto $kY_{h}$ is an isomorphism, this forms a new basis of $C_{i-1}$. Finally, we have $d_{i-1}(kY_{h}^{\prime})=0$ since $d_{i-1}\circ d_{i}=0$, so the resulting basis still respects the stabilizers grow property. Set $D_{i-1}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{h}\oplus\cdots\oplus kY_{n}$ and $K_{i-1}:=kY_{h}^{\prime}$.